OpenStudy (anonymous):

The capacity of a lift is 10 people or 1680 pounds. The capacity will be exceeded if 10 people have weights with a greater mean the 1680/10 pounds. suppose the people have wights that are normally distributed with a mean of 191 pounds and a strandard deviation of 33 lbs. a. find probability that if a person is randomly selected, his weight will be GREATER THAN 168 pounds.

OpenStudy (anonymous):

@CliffSedge check this one out

OpenStudy (anonymous):

Ok, normal distribution calculation. You are finding P(X>168) with N(191,33) You can find the z-score for x=168 and find P(z>z*). Do you know how to do that on the TI-84?

OpenStudy (anonymous):

you might have to remind me..

OpenStudy (anonymous):

the way you start is that you use the equation \[\frac{ given - mean }{ Standard Devation } \] which makes which makes \[\frac{ 168 - 171 }{ 33 } \]

OpenStudy (anonymous):

that equals -.0909090909 and then im supposed to look that up on the z score table which is .4641 and then i subtract that from 1 which equals .5359 but it says the answer is .5362

OpenStudy (anonymous):

171? I thought the mean was 191?

OpenStudy (anonymous):

oops! it's supposed to be 191!!!! sorry

OpenStudy (anonymous):

Well, in your case of using 171, the table value, .5359 that you got is near enough to the more precise value of .5362. You can get the more precise value using the normalcdf function on our calculator.

OpenStudy (anonymous):

it wont work already tried

OpenStudy (anonymous):

Here's a reminder how to do that: Press 2ND-VARS for DISTR menu. Do normalcdf(168,1000,171,33)

OpenStudy (anonymous):

ooooooooh!!!!!!! this whole time i was entering normalcdf(-.0909090909) lol -__-

OpenStudy (anonymous):

why did you use 100 though?

OpenStudy (anonymous):

1000*

OpenStudy (anonymous):

That's an approximation to infinity.

OpenStudy (anonymous):

so what about part b. says find probability that 10 random people selected with have a mean that is greater that 169 pounds?

OpenStudy (anonymous):

oops 168*

OpenStudy (anonymous):

Gah, sorry for the delay; browser crashed . . . Ok, for this one, you will use 1680 as X and 1710 as the mean (since if the mean of ten people is 1680, then the sum of all their weights is 1680)

OpenStudy (anonymous):

You'll need to adjust the standard deviation appropriately as well. Do you know how to do that @hmdp4 ?

OpenStudy (anonymous):

no im not sure how to adjust the sd

OpenStudy (anonymous):

You have to do it in terms of the variance. If the SD is 33, what is the Var?

OpenStudy (anonymous):

i dont remember how to find the variance in this case

OpenStudy (anonymous):

i know it's something like the opposite of the standard devation or something of that matter

OpenStudy (anonymous):

It's simple. Standard deviation = σ; variance = σ^2. Since you're multiplying everything by ten, you multiply the variance by ten, then find the new standard deviation.